### v1

# Emergence of Vorticity and Viscous Stress in Finite Scale Quantum Hydrodynamics

Abstract
The Madelung equations offer a hydrodynamic description of quantum systems, from single particles to quantum fluids. In this formulation, the probability density is mapped onto the fluid density and the phase is treated as a scalar potential generating the velocity field. As examples of potential flows, quantum fluids described in this way are inherently irrotational, but quantum vortices may arise at discrete points where the phase is undefined. In this paper, starting from this irrotational description of a quantum fluid, a coarse-graining procedure is applied to arrive at a macroscopic description of the quantum fluid in which the role of velocity is played by a Favre average of the microscopic velocity field, allowing for finite vorticity at any point in the fluid. It is shown that this vorticity obeys a similar equation to the vorticity equation in classical hydrodynamics and includes a vortex-stretching term. This coarse-graining procedure also gives rise to novel stress terms in the fluid equations, which in the appropriate limit appear analogous to artificial viscous stresses from computational fluid dynamics.
Fluid turbulence is known to be a ubiquitous non-equilibrium phenomenon arising in fluids when the kinetic energy injected at large scales dominates the dissipation of energy at small scales due to viscosityMcComb (1990, 1995); Sreenivasan (1999); Eyink and Sreenivasan (2006). While first identified and extensively studied in classical fluids, turbulence has also been observed in quantum fluids, including both superfluids and ultra-cold atomic gasesVinen and Niemela (2002); Barenghi and Donnelly (2009); Vinen (2010); Paoletti and Lathrop (2011); Tsubota (2014); Barenghi et al. (2014). While there is no consensus on the precise definition of turbulence, it is characterized by spatially complex and temporally aperiodic flow fields and involves processes on many length scales. In three-dimensions, it is well established that turbulence involves a cascade of energy from large to small length scalesKolmogorov (1991); Falkovich (2009); Dubrulle (2019); Carbone and Bragg (2020). While the details of this process are not fully understood, there is significant evidence that, in classical turbulence, this cascade is related to the non-linear phenomena of strain self-amplification and vortex stretchingCarbone and Bragg (2020). In quantum turbulence, this cascade has a remarkably similar behavior at large length scales, with the Kolmogorov spectrum being observed in superfluid 4He and simulations using the Non-linear Schrödinger equation (NLSE)Paoletti and Lathrop (2011). However, in contrast to the case of classical turbulence where the cascade is related to the dynamics of a continuous vorticity field, the cascade in quantum turbulence is caused by the interaction of a discrete number of quantized vortices and transverse modes (Kelvin waves) which propagate along these vorticesBarenghi and Donnelly (2009). While this similarity between classical and quantum turbulence is not unexpected based on intuition from the correspondence principlePaoletti and Lathrop (2011), a full theory establishing the rigorous connection between these two regimes is lacking, due in part to the lack of a single governing equation used to describe quantum turbulence at all length scales.
In contrast to the study of classical turbulence, whose phenomenology is believed to emerge from the dynamics governed by a single equation, the Navier-Stokes equation, current descriptions of quantum turbulence require a hierarchy of modelsBarenghi et al. (2014). The microscopic behavior of the quantum fluid is modeled using the NLSE, or appropriate generalizations to capture strong interactions; the mesoscale behavior is described by vortex-filament models; and macroscopic dynamics are described using the Hall–Vinen–Bekharevich–Khalatnikov (HVBK) modelBarenghi et al. (2014); Zhang et al. (2023). The purpose of the present paper is to demonstrate that a suitable coarse-graining of the microscopic quantum fluid equations does in fact yield fluid equations very similar in form to the Navier-Stokes equation. Specifically, it is shown that the coarse-grained velocity field can possess finite vorticity, which obeys a similar equation to the vorticity equation of classical fluids, and that viscous stress terms also emerge as a consequence of this coarse-graining. The emergence of a novel viscous stress term due to integrating out small scale degrees of freedom has similarities with a recent work which demonstrated a relationship between decoherence and viscosityZhou et al. (2025).
The particular coarse-graining procedure applied in this paper is based on finite scale theory, a method previously studied in the context of classical continuum dynamicsMargolin and Rider (2002); Margolin (2009, 2014, 2019). The motivation for finite scale theory in classical continua is based on the fact that in order to solve many problems in classical continuum dynamics, the continuum equations are almost always approximated by discretizing the domain and forming a mesh, represented by a finite number of points in the original space. In the context of Lagrangian hydrocodes, this discretization amounts to an averaging of the continuum variables over finite size cells. However, these averaged quantities are not necessarily solutions to the original partial differential equations governing the dynamics of the continuum. Finite scale theory provides a connection between the coarse-graining of the continuous domain and the differential equations describing the averaged quantities.
Here I apply a generalized version of finite scale analysis to the Madelung equations of quantum hydrodynamics. I find that, through this coarse-graining procedure new terms emerge at finite length scales, and, moreover, the Favre averaged velocity field describing the evolution of the quantum fluid can possess finite vorticity.
I Madelung Equations
Consider a complex wavefunction $\psi$, whose dynamics are governed by the non-linear Schrödinger equation. Rewriting the complex function in terms of its magnitude and phase, $\psi=\sqrt{\tfrac{1}{m}\rho}\ e^{\tfrac{i}{\hbar}\theta}$, the equations governing the evolution of $\rho$ and $\theta$ may be written as:
$\partial_{t}\rho+\boldsymbol{\nabla}\cdot\left(\rho\textbf{u}\right)$ $=0,$ (1)
$\partial_{t}\textbf{u}+\left(\textbf{u}\cdot\boldsymbol{\nabla}\right)\textbf{u}$ $=\frac{\hbar^{2}}{2m^{2}}\boldsymbol{\nabla}\left(\frac{\nabla^{2}\sqrt{\rho}}{\sqrt{\rho}}\right)-\frac{g}{m}\boldsymbol{\nabla}\rho,$
where we define $\textbf{u}\equiv\boldsymbol{\nabla}\theta/m$. These are the Madelung equationsMadelung (1927) which offer a hydrodynamic interpretation to the wavefunction due to their similarity to the continuity equation and Cauchy momentum equation of fluid mechanics:
$\partial_{t}\rho+\boldsymbol{\nabla}\cdot\left(\rho\textbf{u}\right)$ $=0,$ (2)
$\partial_{t}\textbf{u}+\left(\textbf{u}\cdot\boldsymbol{\nabla}\right)\textbf{u}$ $=\frac{1}{\rho}\boldsymbol{\nabla}\cdot\boldsymbol{\sigma}+\textbf{f},$
where $\rho$ is the mass density of the continuum, u is the velocity field, $\boldsymbol{\sigma}$ is the stress tensor, and f is an external force. When $\boldsymbol{\nabla}\cdot\boldsymbol{\sigma}=-\boldsymbol{\nabla}p$, the above momentum equation becomes the Euler equation, when $\boldsymbol{\nabla}\cdot\boldsymbol{\sigma}=-\boldsymbol{\nabla}p+\mu\nabla^{2}\textbf{u}+\tfrac{1}{3}\mu\boldsymbol{\nabla}\left(\boldsymbol{\nabla}\cdot\textbf{u}\right)$, it becomes the Navier-Stokes equation.
Ignoring the physical interpretations of these equations, it is clear that Eqs. (1) are simply a special case of Eqs. (2), with the choice:
$\frac{1}{\rho}\boldsymbol{\nabla}\cdot\boldsymbol{\sigma}+\textbf{f}=\gamma\boldsymbol{\nabla}\left(\frac{\nabla^{2}\sqrt{\rho}}{\sqrt{\rho}}\right).$
However, it is worth emphasizing that, by definition, the velocity appearing in the Madelung equations is given by the gradient of the phase, $\textbf{u}\equiv\boldsymbol{\nabla}\theta/m$ and therefore the vorticity, $\boldsymbol{\omega}=\boldsymbol{\nabla}\times\textbf{u}$, must vanish except at singular points in the phase. This is not the case in the classical fluid dynamics described by Eqs. (2). In fact, as mentioned above, the vorticity is believed to play an important role in the Kolmogorov cascade in classical turbulence.
II Generalized Finite Scale Analysis
Consider a continuum field $A(\textbf{x},t)$ where x and $t$ are space and time coordinates, respectively. Let us define the coarse-grained field as:
$\langle A(\textbf{x},t)\rangle\equiv\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}d^{3}x^{\prime}f(\textbf{x}^{\prime})A(\textbf{x}+\textbf{x}^{\prime},t),$ (3)
where $f$ is a normalized distribution describing the details of our coarse-graining procedure. The procedure employed in previous finite scale analyses would correspond to the choice of a uniform box distributionMargolin and Rider (2002); Margolin (2009); however, in principle, many other choices could yield similar results. For simplicity, we will assume that $f$ is a separable distribution, $f(\textbf{x})=f_{1}(x_{1})f_{2}(x_{2})f_{3}(x_{3})$ and that each function $f_{i}(x)$ is an even function of the variable $x$. In this section we will derive the general expressions without making further assumptions about the functional form of $f$ and defer a discussion of reasonable choices for $f$ to a later section.
Throughout this work, we denote the moments of this distribution as:
$\mu_{n_{1},n_{2},n_{3}}\equiv\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}d^{3}xf(\textbf{x})x_{1}^{n_{1}}x_{2}^{n_{2}}x_{3}^{n_{3}}.$ (4)
Note that, because $f$ is a separable distribution, these moments are separable: $\mu_{n_{1},n_{2},n_{3}}=\mu_{1,n_{1}}\mu_{2,n_{2}}\mu_{3,n_{3}}$, where $\mu_{i,n}\equiv\int_{-\infty}^{\infty}dxf_{i}(x)x^{n}$. Moreover, because we have assumed that $f$ is even in each of the spatial coordinates, all odd moments are zero, $\mu_{i,2n+1}=0$.
With these conventions, and assuming that $A(\textbf{x},t)$ is sufficiently smooth in the neighborhood of x that it may be expanded in a Taylor series, it is straightforward to derive the following relations:
$\langle A(\textbf{x},t)\rangle$ $=\sum_{n_{1},n_{2},n_{3}=0}^{\infty}\frac{\mu_{n_{1},n_{2},n_{3}}}{n_{1}!n_{2}!n_{3}!}\partial_{x_{1}}^{n_{1}}\partial_{x_{2}}^{n_{2}}\partial_{x_{3}}^{n_{3}}A(\textbf{x},t),$
$\langle\partial_{x_{1}}^{m_{1}}\partial_{x_{2}}^{m_{2}}\partial_{x_{3}}^{m_{3}}A(\textbf{x},t)\rangle$ $=\sum_{n_{1},n_{2},n_{3}=0}^{\infty}\frac{\mu_{n_{1},n_{2},n_{3}}}{n_{1}!n_{2}!n_{3}!}\partial_{x_{1}}^{n_{1}+m_{1}}\partial_{x_{2}}^{n_{2}+m_{2}}\partial_{x_{3}}^{n_{3}+m_{3}}A(\textbf{x},t)=\partial_{x_{1}}^{m_{1}}\partial_{x_{2}}^{m_{2}}\partial_{x_{3}}^{m_{3}}\langle A(\textbf{x},t)\rangle.$
Note that the averaging commutes with spatial derivatives. Now, transform to dimensionless variables, $\xi_{i}=x_{i}/L_{i}$:
$\frac{1}{L_{1}^{m_{1}}L_{2}^{m_{2}}L_{3}^{m_{3}}}\langle\partial^{m_{1}}_{\xi_{1}}\partial^{m_{2}}_{\xi_{2}}\partial^{m_{3}}_{\xi_{3}}A(\boldsymbol{\xi},t)\rangle$ $=\sum_{n_{1},n_{2},n_{3}=0}^{\infty}\frac{1}{n_{1}!n_{2}!n_{3}!}\frac{\mu_{n_{1},n_{2},n_{3}}}{L_{1}^{n_{1}+m_{1}}L_{2}^{n_{2}+m_{2}}L_{3}^{n_{3}+m_{3}}}\partial_{\xi_{1}}^{n_{1}+m_{1}}\partial_{\xi_{2}}^{n_{2}+m_{2}}\partial_{\xi_{3}}^{n_{3}+m_{3}}A(\boldsymbol{\xi},t),$
$\rightarrow\langle\partial^{m_{1}}_{\xi_{1}}\partial^{m_{2}}_{\xi_{2}}\partial^{m_{3}}_{\xi_{3}}A(\boldsymbol{\xi},t)\rangle$ $=\sum_{n_{1},n_{2},n_{3}=0}^{\infty}\frac{1}{n_{1}!n_{2}!n_{3}!}\frac{\mu_{n_{1},n_{2},n_{3}}}{L_{1}^{n_{1}}L_{2}^{n_{2}}L_{3}^{n_{3}}}\partial_{\xi_{1}}^{n_{1}+m_{1}}\partial_{\xi_{2}}^{n_{2}+m_{2}}\partial_{\xi_{3}}^{n_{3}+m_{3}}A(\boldsymbol{\xi},t),$
where $L_{i}$ is a relevant macroscopic length scale. Note that $\mu_{n_{1},n_{2},n_{3}}\sim\ell_{1}^{n_{1}}\ell_{2}^{n_{2}}\ell_{3}^{n_{3}}$, where $\ell_{i}$ are characteristic length scales associated with the spatial dependence of the distribution $f$. We now assume that each of these scales, $\ell_{i}$, is small compared to the corresponding macroscopic scale $L_{i}$:
$\frac{\ell_{i}}{L_{i}}\equiv\eta_{i}<<1.$ (5)
When this condition is satisfied, we may truncate the infinite series above:
$\langle A(\boldsymbol{\xi},t)\rangle$ $=A(\boldsymbol{\xi},t)+\frac{1}{2}\sum_{i=1}^{3}\frac{\mu_{i,2}}{L_{i}^{2}}\partial_{\xi_{i}}^{2}A(\boldsymbol{\xi},t)+\mathcal{O}(\eta_{i}^{4}),$
$\langle\partial_{\xi_{k}}^{2}A(\boldsymbol{\xi},t)\rangle$ $=\partial_{\xi_{k}}^{2}A(\boldsymbol{\xi},t)+\frac{1}{2}\sum_{i=1}^{3}\frac{\mu_{i,2}}{L_{i}^{2}}\partial_{\xi_{i}}^{2}\partial_{\xi_{k}}^{2}A(\boldsymbol{\xi},t)+\mathcal{O}(\eta_{i}^{4}).$
Collecting the terms above and converting back to dimensionful coordinates, we can rewrite the function $A$ in terms of coarse-grained quantities:
$A=\langle A\rangle-\frac{1}{2}\sum_{i=1}^{3}\mu_{i,2}\partial_{x_{i}}^{2}\langle A\rangle+\mathcal{O}(\eta_{i}^{4}),$ (6)
where we omit the explicit dependence of $A$ for brevity.
Following a similar procedure, it is straightforward to derive an expression for the coarse-grained product of two continuous fields $A$ and $B$:
$\langle AB\rangle=\langle A\rangle\langle B\rangle+\sum_{i=1}^{3}\mu_{i,2}\partial_{x_{i}}\langle A\rangle\partial_{x_{i}}\langle B\rangle+\mathcal{O}(\eta_{i}^{4}),$ (7)
and the product of three continuous fields, $A$, $B$, and $C$:
$\langle ABC\rangle$ $=\langle A\rangle\langle B\rangle\langle C\rangle+\sum_{i=1}^{3}\mu_{i,2}\left[\partial_{x_{i}}\langle A\rangle\partial_{x_{i}}\langle B\rangle\langle C\rangle\right$. (8)
$\left.+\partial_{x_{i}}\langle A\rangle\langle B\rangle\partial_{x_{i}}\langle C\rangle+\langle A\rangle\partial_{x_{i}}\langle B\rangle\partial_{x_{i}}\langle C\rangle\right]$
$+\mathcal{O}(\eta_{i}^{4}).$
Eq. (7) represents a generalization of Eq (2.3) from Reference [(18)] to a coarse-graining of the form shown in Eq. (3). These relations simplify somewhat if we assume an isotropic coarse-graining, so that $\mu_{1,n}=\mu_{2,n}=\mu_{3,n}=\mu_{n}$. Defining the microscopic length scale $\ell\equiv\sqrt{\mu_{2}}$ and taking the macroscopic length scale $L$ to be much larger than $\ell$, we can define the small parameter $\eta\equiv\ell/L$. With these definitions we find that the above relations become:
$A$ $=\langle A\rangle-\frac{\ell^{2}}{2}\nabla^{2}\langle A\rangle+\mathcal{O}(\eta^{4}),$ (9)
$\langle AB\rangle$ $=\langle A\rangle\langle B\rangle+\ell^{2}\boldsymbol{\nabla}\left\langle A\right\rangle\cdot\boldsymbol{\nabla}\left\langle B\right\rangle+\mathcal{O}(\eta^{4}),$
$\langle ABC\rangle$ $=\left\langle A\right\rangle\left\langle B\right\rangle\left\langle C\right\rangle+\ell^{2}\left[\boldsymbol{\nabla}\left\langle A\right\rangle\cdot\boldsymbol{\nabla}\left\langle B\right\rangle\left\langle C\right\rangle\right$.
$\left.+\boldsymbol{\nabla}\left\langle A\right\rangle\left\langle B\right\rangle\cdot\boldsymbol{\nabla}\left\langle C\right\rangle+\left\langle A\right\rangle\boldsymbol{\nabla}\left\langle B\right\rangle\cdot\boldsymbol{\nabla}\left\langle C\right\rangle\right]$
$+\mathcal{O}(\eta^{4}).$
We have now obtained the results we need to coarse-grain the Madelung equations. In all that follows we assume an isotropic finite scale coarse-graining with the conventions defined above.
III Finite Scale Madelung Equations
III.1 Continuity Equation
Using the averaging defined in Eq. (3), we can write the continuity equation as:
$\partial_{t}\langle\rho\rangle+\boldsymbol{\nabla}\cdot\langle\rho\textbf{u}\rangle$ $=0.$
Defining the Favre average:
$\langle\langle\textbf{u}\rangle\rangle=\frac{\langle\rho\textbf{u}\rangle}{\langle\rho\rangle},$ (10)
we find that the coarse-grained continuity equation can be written as:
$\partial_{t}\langle\rho\rangle+\boldsymbol{\nabla}\cdot\left(\langle\rho\rangle\langle\langle\textbf{u}\rangle\rangle\right)=0.$ (11)
Therefore, we see that the mass continuity equation applies exactly to these averaged quantities, where the Favre averaged velocity plays the role of a macroscopic velocity field. We can expand this Favre averaged velocity in terms of the finite scale averaged velocity using Eq. (7) to obtain:
$\langle\langle\textbf{u}\rangle\rangle=\left\langle\textbf{u}\right\rangle+\frac{\ell^{2}}{\langle\rho\rangle}\boldsymbol{\nabla}\left\langle\rho\right\rangle\cdot\boldsymbol{\nabla}\left\langle\textbf{u}\right\rangle+\mathcal{O}(\eta^{4}).$ (12)
These two averages are, in general, different.
III.2 Momentum Equation
We now turn our attention to the momentum equation in Eqs. (1). Multiplying the microscopic version of the momentum equation by $\rho$, we find:
$\partial_{t}\left(\rho\textbf{u}\right)+\sum_{i=1}^{3}\partial_{x_{i}}\left(u_{i}\rho\textbf{u}\right)=\frac{\hbar^{2}}{2m^{2}}\rho\boldsymbol{\nabla}\left(\frac{\nabla^{2}\sqrt{\rho}}{\sqrt{\rho}}\right)-\frac{g}{m}\rho\boldsymbol{\nabla}\rho,$ (13)
where we have used the microscopic continuity equation to simplify the left hand side. We will now apply the finite scale coarse-graining procedure to Eq. (13). Due to the complexity of some of these terms, more care is required in applying the procedure to this equation than in the case of Eq. (11).
The first term on the left hand side may be written exactly as:
$\left\langle\partial_{t}\left(\rho\textbf{u}\right)\right\rangle$ $=\partial_{t}\left(\left\langle\rho\right\rangle\frac{\left\langle\rho\textbf{u}\right\rangle}{\left\langle\rho\right\rangle}\right)=\partial_{t}\left(\left\langle\rho\right\rangle\langle\langle\textbf{u}\rangle\rangle\right),$
$=\langle\langle\textbf{u}\rangle\rangle\partial_{t}\left\langle\rho\right\rangle+\left\langle\rho\right\rangle\partial_{t}\langle\langle\textbf{u}\rangle\rangle,$
$=-\langle\langle\textbf{u}\rangle\rangle\left[\langle\langle\textbf{u}\rangle\rangle\cdot\boldsymbol{\nabla}\left\langle\rho\right\rangle+\left\langle\rho\right\rangle\boldsymbol{\nabla}\cdot\langle\langle\textbf{u}\rangle\rangle\right]$
$+\left\langle\rho\right\rangle\partial_{t}\langle\langle\textbf{u}\rangle\rangle,$
$=-\sum_{i=1}^{3}\partial_{x_{i}}\left[\left\langle\left\langle u_{i}\right\rangle\right\rangle\left\langle\rho\right\rangle\langle\langle\textbf{u}\rangle\rangle\right]$
$+\left\langle\rho\right\rangle\left[\partial_{t}\langle\langle\textbf{u}\rangle\rangle+\langle\langle\textbf{u}\rangle\rangle\cdot\boldsymbol{\nabla}\ \langle\langle\textbf{u}\rangle\rangle\right],$
where we have used Eq. (11) to eliminate the time derivative of $\left\langle\rho\right\rangle$. Collecting terms and dividing by $\left\langle\rho\right\rangle$ we find the exact momentum equation relating the coarse-grained quantities:
$\partial_{t}\langle\langle\textbf{u}\rangle\rangle+\langle\langle\textbf{u}\rangle\rangle\cdot\boldsymbol{\nabla}\ \langle\langle\textbf{u}\rangle\rangle$ $=\frac{\hbar^{2}}{2m^{2}}\left\langle\left\langle\boldsymbol{\nabla}\left(\frac{\nabla^{2}\sqrt{\rho}}{\sqrt{\rho}}\right)\right\rangle\right\rangle-\frac{g}{m}\left\langle\left\langle\boldsymbol{\nabla}\rho\right\rangle\right\rangle+\frac{1}{\langle\rho\rangle}\sum_{i=1}^{3}\left\{\partial_{x_{i}}\left[\left\langle\left\langle u_{i}\right\rangle\right\rangle\left\langle\rho\right\rangle\langle\langle\textbf{u}\rangle\rangle-\left\langle u_{i}\rho\textbf{u}\right\rangle\right]\right\}.$ (14)
Notice that the left hand side is of the same form as in the original Madelung equations, but with u replaced with $\langle\langle\textbf{u}\rangle\rangle$. Notice, also, the appearance of three terms on the right hand side of the equation: a quantum stress term:
$\textbf{S}_{Q}=\frac{\hbar^{2}}{2m^{2}}\left\langle\left\langle\boldsymbol{\nabla}\left(\frac{\nabla^{2}\sqrt{\rho}}{\sqrt{\rho}}\right)\right\rangle\right\rangle,$ (15)
a stress term due to microscopic interactions:
$\textbf{S}_{I}=\frac{g}{m}\left\langle\left\langle\boldsymbol{\nabla}\rho\right\rangle\right\rangle,$ (16)
and a classical stress term:
$\textbf{S}_{C}=\frac{1}{\langle\rho\rangle}\sum_{i=1}^{3}\left\{\partial_{x_{i}}\left[\left\langle\left\langle u_{i}\right\rangle\right\rangle\left\langle\rho\right\rangle\langle\langle\textbf{u}\rangle\rangle-\left\langle u_{i}\rho\textbf{u}\right\rangle\right]\right\}.$ (17)
The quantum stress is simply the Favre average of the quantum pressure term in the Madelung equations, the interaction stress is, similarly, the Favre average of the term due to interactions. The classical stress term has the form of the divergence of a stress tensor. In principle, one could perform the averages in Eq. (17) for specific quantum systems to study how these terms depend on the macroscopic flow variables. However, such a study is beyond the scope of the current work. Instead, following the finite scale analysis for classical fluids, we will expand each term in Eq. (17) perturbatively in $\ell^{2}$.
Expanding the average triple product in the classical stress term, we find:
$\left\langle u_{i}\rho\textbf{u}\right\rangle$ $=\langle u_{i}\rangle\langle\rho\rangle\langle\textbf{u}\rangle+\ell^{2}\left[\boldsymbol{\nabla}\langle u_{i}\rangle\cdot\boldsymbol{\nabla}\langle\rho\rangle\langle\textbf{u}\rangle\right$.
$\left.+\boldsymbol{\nabla}\langle u_{i}\rangle\langle\rho\rangle\cdot\boldsymbol{\nabla}\ \langle\textbf{u}\rangle+\langle u_{i}\rangle\boldsymbol{\nabla}\langle\rho\rangle\cdot\boldsymbol{\nabla}\ \langle\textbf{u}\rangle\right]+\mathcal{O}(\eta^{4}),$
where we have simply used the identity in Eq. (8). Similarly, the triple product of averages may be written as:
$\left\langle\left\langle u_{i}\right\rangle\right\rangle\left\langle\rho\right\rangle\langle\langle\textbf{u}\rangle\rangle$ $=\left\langle u_{i}\right\rangle\left\langle\rho\right\rangle\left\langle\textbf{u}\right\rangle+\ell^{2}\left[\left\langle u_{i}\right\rangle\boldsymbol{\nabla}\left\langle\rho\right\rangle\cdot\boldsymbol{\nabla}\left\langle\textbf{u}\right\rangle\right$.
$\left.+\boldsymbol{\nabla}\left\langle\rho\right\rangle\cdot\boldsymbol{\nabla}\left\langle u_{i}\right\rangle\left\langle\textbf{u}\right\rangle\right]+\mathcal{O}(\eta^{4}).$
Combining these results, we find that Eq. (17) may be written as:
$\textbf{S}_{C}$ $=-\frac{1}{\langle\rho\rangle}\ell^{2}\sum_{i=1}^{3}\partial_{x_{i}}\left[\left\langle\rho\right\rangle\boldsymbol{\nabla}\left\langle\left\langle u_{i}\right\rangle\right\rangle\cdot\boldsymbol{\nabla}\ \langle\langle\textbf{u}\rangle\rangle\right]+\mathcal{O}(\eta^{4}),$
where we have replaced $\left\langle\textbf{u}\right\rangle$ with $\langle\langle\textbf{u}\rangle\rangle$ because they are equivalent at this order in $\ell^{2}$. Notice that, when expanded to leading order in $\eta=\ell/L$, the classical stress appears to be reminiscent of an artificial viscous stress term from computational fluid dynamics.
To gain insight into the relative importance of each of the stress terms on the right hand side, we define a characteristic magnitude of the velocity field $U_{0}$ and a characteristic macroscopic density $\rho_{0}$. Combining the characteristic velocity with the macroscopic length scale $L$ we define the time scale $\tau=L/U_{0}$. With these scales defined, it is straightforward to arrive at the non-dimensionalized momentum equation:
$\partial_{\tau}\tilde{\textbf{u}}+\tilde{\textbf{u}}\cdot\boldsymbol{\nabla}_{\boldsymbol{\xi}}\ \tilde{\textbf{u}}$ $=\frac{\hbar^{2}}{2m^{2}L^{2}U_{0}^{2}}\tilde{\textbf{S}}_{Q}-\frac{g\rho_{0}}{mU_{0}^{2}}\tilde{\textbf{S}}_{I}-\frac{\ell^{2}}{L^{2}}\tilde{\textbf{S}}_{C},$ (18)
where we define the normalized stress terms:
$\tilde{\textbf{S}}_{Q}=L^{3}\left\langle\left\langle\boldsymbol{\nabla}\left(\frac{\nabla^{2}\sqrt{\rho}}{\sqrt{\rho}}\right)\right\rangle\right\rangle,$ (19)
$\tilde{\textbf{S}}_{I}=\frac{L}{\rho_{0}}\left\langle\left\langle\boldsymbol{\nabla}\rho\right\rangle\right\rangle,$ (20)
and:
$\tilde{\textbf{S}}_{C}=\frac{L^{2}}{\langle\rho\rangle U_{0}^{2}}\sum_{i=1}^{3}\partial_{x_{i}}\left[\left\langle\rho\right\rangle\boldsymbol{\nabla}\left\langle\left\langle u_{i}\right\rangle\right\rangle\cdot\boldsymbol{\nabla}\ \langle\langle\textbf{u}\rangle\rangle\right].$ (21)
We now see that the relative importance of each stress term is determined, in part, by the magnitudes of the three dimensionless parameters:
$q$ $=\frac{\hbar^{2}}{2m^{2}L^{2}U_{0}^{2}},$ (22)
$\gamma$ $=\frac{g\rho_{0}}{mU_{0}^{2}},$
$\eta$ $=\frac{\ell^{2}}{L^{2}}.$
From these quantities we can see that classical and quantum stresses are comparable in magnitude when the coarse-graining scale is chosen to be:
$\ell\sim\ell_{Q}=\frac{\hbar}{\sqrt{2}mU_{0}}.$ (23)
This sets a natural finite scale at which quasi-classical behavior should be expected.
For a fluid with a macroscopic velocity of the order $U_{0}\sim 1$m/s, and $m\sim m_{p}=1.67262192\times 10^{-27}$ kg, we find this coarse-graining scale to be $\ell_{Q}\sim 45$ nm. From this analysis, we conclude that when the interaction energy density is sufficiently low compared to the macroscopic kinetic energy, $U_{0}^{2}/\rho_{0}>>g/m$, and we are only concerned with behavior of the flow fields on scales such that $\ell>>\frac{\hbar}{\sqrt{2}mU_{0}}$, then the dynamics of this system may be described by a set of classical fluid equations.
Another possible choice for $U_{0}$ is the characteristic particle velocity for an ideal gas with temperature $T$: $U_{0}=\sqrt{3k_{B}T/m}$. Inserting this into Eq. (23) we find that the quasi-classical coarse-graining scale becomes: $\ell_{Q}=\sqrt{\frac{\hbar^{2}}{6mk_{B}T}}$, which is roughly a factor of 6 smaller than the thermal de Broglie wavelength: $\lambda=\sqrt{\frac{2\pi\hbar^{2}}{mk_{B}T}}$. Therefore, choosing $\ell>>\lambda$ is a sufficient condition for classical behavior, when the interactions can be neglected, as one would expect from the correspondence principle.
III.3 Vorticity Equation
While the velocity appearing in the microscopic Madelung equations, Eq. (1), is by definition irrotational, $\boldsymbol{\nabla}\times\textbf{u}=0$, the same is not necessarily true for the Favre averaged velocity appearing in the finite scale Madelung equations, Eqs. (11) and (14). In general, the vorticity of this coarse-grained field has the form:
$\boldsymbol{\omega}\equiv\boldsymbol{\nabla}\times\langle\langle\textbf{u}\rangle\rangle$ $=\boldsymbol{\nabla}\times\frac{\left\langle\rho\textbf{u}\right\rangle}{\left\langle\rho\right\rangle},$
$=\frac{\left\langle\boldsymbol{\nabla}\rho\times\textbf{u}\right\rangle}{\left\langle\rho\right\rangle}-\frac{\boldsymbol{\nabla}\left\langle\rho\right\rangle\times\left\langle\rho\textbf{u}\right\rangle}{\left\langle\rho\right\rangle^{2}},$
$=\frac{1}{\left\langle\rho\right\rangle}\left[\left\langle\boldsymbol{\nabla}\rho\times\textbf{u}\right\rangle-\boldsymbol{\nabla}\left\langle\rho\right\rangle\times\langle\langle\textbf{u}\rangle\rangle\right].$
This expression is exact and applies to the coarse-grained velocity field at all orders in the finite scale, $\ell$. There is no reason to expect that it should necessarily be zero, in general. Moreover, we can expand the right hand side to leading order in $\ell^{2}$ to arrive at:
$\omega_{i}$ $=\frac{\ell^{2}}{\left\langle\rho\right\rangle^{2}}\sum_{j,k,l}\epsilon_{ijk}\left[\left\langle\rho\right\rangle\partial_{x_{j}}\partial_{x_{l}}\left\langle\rho\right\rangle\right$. (24)
$\left.-\partial_{x_{j}}\left\langle\rho\right\rangle\ \partial_{x_{l}}\left\langle\rho\right\rangle\right]\partial_{x_{l}}\left\langle u_{k}\right\rangle+\mathcal{O}(\eta^{4}).$
In the next section we will consider a concrete example of a solution to the Madelung equations whose coarse-grained description possesses finite vorticity, thus demonstrating that vorticity is present in the finite scale description of even some irrotational flows. Before we construct that example, we will consider the equation describing the dynamics of the finite scale vorticity field. This can be obtained, in the same way as is done in classical hydrodynamics, by taking the curl of the momentum equation, Eq. (14):
$\partial_{t}\boldsymbol{\omega}$ $+\langle\langle\textbf{u}\rangle\rangle\cdot\boldsymbol{\nabla}\ \boldsymbol{\omega}=\boldsymbol{\omega}\cdot\boldsymbol{\nabla}\ \langle\langle\textbf{u}\rangle\rangle-\boldsymbol{\omega}\ \boldsymbol{\nabla}\cdot\langle\langle\textbf{u}\rangle\rangle$ (25)
$+\boldsymbol{\nabla}\times\textbf{S}_{Q}-\boldsymbol{\nabla}\times\textbf{S}_{I}+\boldsymbol{\nabla}\times\textbf{S}_{C}.$
Note that, just as in classical hydrodynamics, the left hand side is the Lagrangian derivative of the vorticity field, that is the time derivative of the vorticity in the frame of the fluid. Also, just as in classical hydrodynamics, a vortex-stretching term, $\boldsymbol{\omega}\cdot\boldsymbol{\nabla}\ \langle\langle\textbf{u}\rangle\rangle$, emerges on the right hand side, implying that vorticity will, generally, be enhanced when the velocity field increases along the direction parallel to $\boldsymbol{\omega}$. This phenomenon is thought to play an important role in the cascade of energy from larger to smaller length scales in classical turbulence. Since it arises naturally upon coarse-graining, it stands to reason that similar phenomena should be expected to play a role in macroscopic effective descriptions of quantum turbulence as well, even though vorticity is, strictly-speaking, absent in quantum fluids.
Simplifications can be made if we expand this equation to leading order in the dimensionless parameters in Eq. (22), $q$, $\gamma$, and $\eta$. In this case, to leading order $\boldsymbol{\nabla}\times\textbf{S}_{Q}=\boldsymbol{\nabla}\times\textbf{S}_{I}=0$, so that:
$\partial_{t}\boldsymbol{\omega}$ $+\langle\langle\textbf{u}\rangle\rangle\cdot\boldsymbol{\nabla}\ \boldsymbol{\omega}=\boldsymbol{\omega}\cdot\boldsymbol{\nabla}\ \langle\langle\textbf{u}\rangle\rangle-\boldsymbol{\omega}\ \boldsymbol{\nabla}\cdot\langle\langle\textbf{u}\rangle\rangle$ (26)
$-\ell^{2}\boldsymbol{\nabla}\times\left\{\frac{1}{\langle\rho\rangle}\sum_{i=1}^{3}\partial_{x_{i}}\left[\left\langle\rho\right\rangle\boldsymbol{\nabla}\left\langle\left\langle u_{i}\right\rangle\right\rangle\cdot\boldsymbol{\nabla}\ \langle\langle\textbf{u}\rangle\rangle\right]\right\}$
$+\mathcal{O}(\eta^{4}).$
From this equation we see that at leading order in these parameters the evolution of the vorticity only depends on coarse-grained quantities and does not depend explicitly on any of the parameters appearing in the original microscopic description.
IV Example Flow Field: A Line Vortex
To better understand the emergence of vorticity in the finite scale theory outlined in the previous section, we will now consider an example flow field which is chosen to be amenable to a semi-analytic treatment while also exhibiting emergent vorticity when coarse-grained.
With these motivations in mind, we choose the velocity field to be that associated with a classical line vortex and a density given by a Gaussian distribution in three dimensions. The wavefunction associated with these choices is given by:
$\psi_{0}=\frac{1}{(2\pi)^{3/4}\sigma^{3/2}}e^{-\frac{\textbf{x}\cdot\textbf{x}}{4\sigma^{2}}}e^{i\frac{m}{\hbar}\frac{\Gamma}{2\pi}\tan^{-1}\left(\frac{x_{2}}{x_{1}}\right)}.$ (27)
We can see that $\psi$ is normalized in space, since we have chosen $\psi_{0}^{*}\psi_{0}$ to be a normalized Gaussian distribution. While it is not a solution of the Schrödinger equation, it is a valid initial wavefunction which would evolve in time according to the Schrödinger equation and it is a relatively simple example of a wavefunction whose coarse-grained hydrodynamic description can be shown to possess finite vorticity.
With this choice of wavefunction, the initial density and velocity fields are:
$\rho$ $=\frac{m}{(2\pi)^{3/2}\sigma^{3}}e^{-\frac{\textbf{x}\cdot\textbf{x}}{2\sigma^{2}}},$ (28)
u $=\frac{\Gamma}{2\pi}\frac{1}{x_{1}^{2}+x_{2}^{2}}\left[-x_{2}\hat{x}_{1}+x_{1}\hat{x}_{2}\right],$
where $\Gamma$ is the circulation around the line vortex. We note that, while this flow field does possess finite circulation, given by $\Gamma$, it is actually irrotational in that $\boldsymbol{\nabla}\times\textbf{u}=0$ everywhere, except at the singular point $x_{1}=x_{2}=0$, as one can easily verify from the expression above.
To perform the coarse-graining procedure discussed in the previous section, we must choose a distribution to use as a filter. For convenience, we choose the Gaussian distribution:
$f(\textbf{x})=\frac{1}{(2\pi)^{3/2}\ell^{3}}e^{-\frac{\textbf{x}\cdot\textbf{x}}{2\ell^{2}}}.$ (29)
This form has a few advantages, the most important one for the current task is that the product $\rho f$ is fairly easy to integrate. In principle, other choices of $f$ may be more or less convenient for understanding other problems, but a complete characterization of this topic is beyond the scope of the current work.
We can now evaluate the coarse-grained fields, $\left\langle\rho\right\rangle$ and $\langle\langle\textbf{u}\rangle\rangle$. Starting with the density, we find:
$\left\langle\rho\right\rangle$ $=\frac{m\prod_{i=1}^{3}\int_{-\infty}^{\infty}dx^{\prime}_{i}\exp\left[-\frac{\sigma^{2}x_{i}^{\prime 2}+\ell^{2}\left(x_{i}^{\prime}+x_{i}\right)^{2}}{2\sigma^{2}\ell^{2}}\right]}{(2\pi)^{3}\sigma^{3}\ell^{3}},$
$=\frac{m\exp\left\{-\frac{\textbf{x}\cdot\textbf{x}}{2\left(\sigma^{2}+\ell^{2}\right)}\right\}}{(2\pi)^{3}}\left(\frac{2}{\sigma^{2}+\ell^{2}}\right)^{3/2}$
$\times\prod_{i=1}^{3}\int_{-\infty}^{\infty}du_{i}\exp\left[-\left(u_{i}+\frac{\ell x_{i}}{\sigma\sqrt{2\left(\sigma^{2}+\ell^{2}\right)}}\right)^{2}\right],$
performing the final Gaussian integrals we have:
$\left\langle\rho\right\rangle=\frac{m\exp\left\{-\frac{\textbf{x}\cdot\textbf{x}}{2\left(\sigma^{2}+\ell^{2}\right)}\right\}}{\left[2\pi\left(\sigma^{2}+\ell^{2}\right)\right]^{3/2}}.$ (30)
As expected, in the limit of $\ell\rightarrow 0$ this equals $\rho$.
This allows us to write the velocity field as:
$\left\langle\left\langle u_{1}\right\rangle\right\rangle$ $=-\frac{\Gamma\sqrt{\ell^{-2}+\sigma^{-2}}}{2^{3/2}\pi^{2}}\int_{-\infty}^{\infty}d\zeta_{1}\int_{-\infty}^{\infty}d\zeta_{2}$ (31)
$\times\zeta_{2}\frac{\exp\left[-\left(\zeta_{1}-\frac{x_{1}}{\alpha}\right)^{2}-\left(\zeta_{2}-\frac{x_{2}}{\alpha}\right)^{2}\right]}{\zeta_{1}^{2}+\zeta_{2}^{2}},$
$\left\langle\left\langle u_{2}\right\rangle\right\rangle$ $=\frac{\Gamma\sqrt{\ell^{-2}+\sigma^{-2}}}{2^{3/2}\pi^{2}}\int_{-\infty}^{\infty}d\zeta_{1}\int_{-\infty}^{\infty}d\zeta_{2}$
$\times\zeta_{1}\frac{\exp\left[-\left(\zeta_{1}-\frac{x_{1}}{\alpha}\right)^{2}-\left(\zeta_{2}-\frac{x_{2}}{\alpha}\right)^{2}\right]}{\zeta_{1}^{2}+\zeta_{2}^{2}},$
$\left\langle\left\langle u_{3}\right\rangle\right\rangle$ $=0,$
where we define the constant $\alpha\equiv\ell\sqrt{2\left(\sigma^{2}+\ell^{2}\right)}/\sigma$. Since the purpose of this example is to demonstrate the presence of finite vorticity, we will compute $\boldsymbol{\omega}$ directly without explicitly evaluating $\langle\langle\textbf{u}\rangle\rangle$. This can be done by taking the curl of the velocity defined above:
$\boldsymbol{\omega}$ $=\boldsymbol{\nabla}\times\langle\langle\textbf{u}\rangle\rangle=\hat{x}_{3}\left[\partial_{x_{1}}\left\langle\left\langle u_{2}\right\rangle\right\rangle-\partial_{x_{2}}\left\langle\left\langle u_{1}\right\rangle\right\rangle\right]=\hat{x}_{3}\omega,$
where $\omega$ is given by the integral:
$\omega(x_{1},x_{2})$ $=\frac{\Gamma}{2\pi^{2}\ell^{2}}\left\{\pi-\int_{-\infty}^{\infty}d\zeta_{1}\int_{-\infty}^{\infty}d\zeta_{2}\right$. (32)
$\left.\times\frac{\zeta_{1}\frac{x_{1}}{\alpha}+\zeta_{2}\frac{x_{2}}{\alpha}}{\zeta_{1}^{2}+\zeta_{2}^{2}}e^{-\left(\zeta_{1}-\tfrac{x_{1}}{\alpha}\right)^{2}-\left(\zeta_{2}-\tfrac{x_{2}}{\alpha}\right)^{2}}\right\}.$
Defining the radius from the vortex core as $r=\sqrt{x_{1}^{2}+x_{2}^{2}}$ we can write the vorticity just as a function of $r$:
$\omega(r)$ $=\frac{\Gamma}{2\pi^{2}\ell^{2}}\left\{\pi-\frac{r}{\alpha}\int_{0}^{\infty}dr^{\prime}\int_{0}^{2\pi}d\phi\right$. (33)
$\left.\times e^{-r^{\prime 2}+\frac{2r}{\alpha}r^{\prime}\cos\phi}e^{-\frac{r^{2}}{\alpha^{2}}}\cos\phi\right\},$
$=\frac{\Gamma}{2\pi\ell^{2}}\left\{1-\frac{1}{2\sqrt{\pi}}\frac{r}{\alpha}\int_{0}^{2\pi}d\phi\right$.
$\left.\times e^{-\frac{r^{2}}{\alpha^{2}}\sin^{2}\phi}\cos\phi\ \text{erf}\left(\frac{r\cos\phi}{\alpha}\right)\right\}.$
Written in this form, it is clear that, along the center line of the vortex $r=0$, the vorticity is equal to the vortex strength divided by the area $\ell^{2}$:
$\omega_{0}=\frac{\Gamma}{2\pi\ell^{2}}.$ (34)
Figure 1: Plot of emergent vorticity as given by Eq. (33) as a function of radius from vortex center. Vorticity is normalized to $\omega_{0}$ as defined in Eq. (34) and radii are expressed in units of $\alpha=\ell\sqrt{2\left(\sigma^{2}+\ell^{2}\right)}/\sigma$.
Further insight can be gained by expanding the error function using its Taylor series representation:
$\omega(r)$ $=\omega_{0}\left\{1-\frac{1}{\pi}\sum_{n=0}^{\infty}\left(\frac{r}{\alpha}\right)^{2(n+1)}\int_{0}^{2\pi}d\phi\right$.
$\left.\times e^{-\frac{r^{2}}{\alpha^{2}}\sin^{2}\phi}\left(\cos\phi\right)^{2(n+1)}\frac{\left(-1\right)^{n}}{n!\left(2n+1\right)}\right\}.$
Taking the leading order term in $r/\alpha$ we find that for $r<<\alpha=\ell\sqrt{2\left(\sigma^{2}+\ell^{2}\right)}/\sigma$ the vorticity is given approximately by:
$\omega(r)=\omega_{0}\left[1-\frac{1}{2}\frac{r^{2}\sigma^{2}}{\ell^{2}\left(\sigma^{2}+\ell^{2}\right)}\right].$ (35)
In the opposite limit, $r>>\alpha$ we note that $\frac{r}{\alpha}e^{-\frac{r^{2}}{\alpha^{2}}\sin^{2}\phi}\approx\sqrt{\pi}\delta\left(\sin\phi\right)$, so that:
$\lim_{r\rightarrow\infty}\omega(r)$ $=\omega_{0}\left\{1-\frac{1}{2}\int_{0}^{2\pi}d\phi\cos\phi\right$. (36)
$\left.\times\left[\delta\left(\phi\right)+\delta\left(\phi-\pi\right)\right]\lim_{r\rightarrow\infty}\text{erf}\left(\frac{r\cos\phi}{\alpha}\right)\right\},$
$=0.$
This demonstrates that the finite scale coarse graining procedure can indeed give rise to finite vorticity, even though the underlying microscopic flow is irrotational. This vorticity is peaked at the vortex core and decreases to zero at large radii compared to $\alpha=\ell\sqrt{2\left(\sigma^{2}+\ell^{2}\right)}/\sigma$, see Figure 1. Furthermore, we note that in the limit of $\ell\rightarrow 0$ the peak at the vortex core diverges while the value at any finite radius away from the core vanishes, in agreement with the singular behavior of the microscopic flow field. In this sense, we can think of the coarse graining procedure as smoothing out the singular behavior of the line vortex and distributing the circulation into a region with finite vorticity.
V Discussion and Conclusions
In this paper we have derived a set of hydrodynamic equations for spatially coarse-grained quantum systems. The coarse-graining procedure we employed represents a generalization of the finite scale theory which has been previously applied in the study of classical fluids. As in the case of classical fluids, we find that the appropriate velocity field in the macroscopic fluid equations is a Favre averaged velocity field and that the coarse-graining gives rise to emergent viscous stresses in the momentum equations. We have also shown that this macroscopic velocity field can possess an emergent vorticity, even though the microscopic velocity is irrotational. The equation governing the evolution of the vorticity was also derived and shown to possess the same vortex-stretching term that is present in classical fluids. Furthermore, expanding the stress terms to leading order in small parameters, we found that the equation governing the vorticity evolution depends only on coarse-grained quantities and has no dependence on the parameters describing the microscopic quantum model, consistent with its emergence at finite scales.
The existence of finite vorticity was demonstrated explicitly in Sec.IV using the specific example of an irrotational line vortex. It was shown that the emergent vorticity in that example is proportional to the vortex strength and that the coarse-graining acts to smooth out the singular behavior of the underlying microscopic velocity field and distribute finite vorticity through the fluid around the vortex. It should be noted that, in that case, the small parameter associated with the coarse-graining length scale appeared in the denominator of the vorticity due to singular behavior in the microscopic theory. This implies that for microscopic fields with singularities, the finite scale expansion presented in this work may require modifications. However, while the small parameter expansion may fail for singular microscopic fields, the exact expressions presented in this work should still apply and the general conclusions about the emergence of vorticity and novel viscous stress terms in the hydrodynamic equations remain valid even for singular fields.
This analysis offers a novel perspective on quantum fluids and the emergence of classical fluid dynamics from microscopic processes. It suggests that the main terms contributing to vorticity dynamics at large scales are, in fact, universal, in that they do not depend on the microscopic degrees of freedom. This has important implications because vorticity dynamics play a major role in the energy cascade from large to small length scales in classical turbulence.
Acknowledgements.
I would like to thank Alexander Balatsky, Roy Baty, and Len Margolin for insightful discussions and comments on various topics related to fluid dynamics, both quantum and classical.